Express each number as a product of its prime factors:
7429

Grade:12

1 Answers

Harshit Singh
3 years ago
Dear Student

7429
By Taking the LCM of 7429, we will get the product of its prime factor.
Hence, 7429 = 17 × 19 × 23 × 1 = 17 × 19 × 23

​Thanks